a cylindrical silo has a hemispherical top. the cylinder has a height of 40 m and a base diameter of 12m.

b. the silo should be filled no more than 80% capacity. how much grain can be put in the silo>

c. a truck with a bin measuring 8 m by 4 m by 3.5m delivers grain. how many truckloads would fill a silo to its recommended capacity
1
Expert's answer
2020-12-10T20:13:21-0500

Consider the cylindrical silo with hemispherical top as shown in the figure below:





The total volume of the silo is,


"V=\pi (6^2)(40)+\frac{2}{3} \pi (6^3)"


"=1440\pi+144\pi"


"=1584\pi"


The amount of grain put into the silo is,


"A=1584\pi(\frac{80}{100})"


"=1267.20\pi"


The number of truckloads to fill to silo to its recommended capacity of 80% is,


"n=\frac{1267.20\pi}{(8)(4)(3.5)}"


"=\frac{396}{35}\pi"


"\approx 36"


Therefore, the number of truckloads to fill the silo to its recommended capacity is "n\approx36"